from the deck of 52 playing cards how many ways can be drawn such that:
a. they both red?
b. one is a diamond card and the other is a spade card?

Answer by stanbon(75887)

from the deck of 52 playing cards how many ways can be drawn such that:
a. they both red?
Ans: 26C2 = (26*25)/(1*2) = 325 ways
-------------------------------------------
b. one is a diamond card and the other is a spade card?
Ans: 13*13 = 169 ways
==========================
Cheers,
Stan H.